OpenStudy (anonymous):

Solve for x: 3x - 5 = 2x + 6

OpenStudy (anonymous):

3x - 5 = 2x + 6 +5 3x = 2x + 11 -2x x =11

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem! Glad to have helped. xD